A university has 10,000 students of which 35% are male and 65% are female. If a class of 30 students is chosen at random from the university population, find the mean and variance of the number of male students.
Answer by stanbon(75887)

A university has 10,000 students of which 35% are male and 65% are female. If a class of 30 students is chosen at random from the university population, find the mean and variance of the number of male students.
------
Binomial:
mean = np = 10,000*0.35 = 3500
---
variance = npq = 3500*0.65 = 2275
==================================
Cheers,
Stan H.
===========